A body starting from rest travels for 100s with uniform acceleration of 1.5m/s . What distance does it cover in the last 2 seconds


1
Expert's answer
2021-03-24T19:43:14-0400

"\Delta s=s_{100}-s_{98}=\frac{a}{2}(t_{100}^2-t_{98}^2)=\frac{1.5}{2}\cdot(100^2-98^2)=297~m."